OpenStudy (anonymous):

A system of equations is shown below: 2x = −y + 6 −4x + 3y = 8 What is the solution to this system of equations? (−1, −4) (1, 4) (4, 1) (−4, −1)

OpenStudy (anonymous):

@imqwerty

imqwerty (imqwerty):

try doing this->multiply the 1st equation by 2 and add both of them

jimthompson5910 (jim_thompson5910):

You can also graph the two equations. The point of intersection is the solution to the system https://www.desmos.com/calculator/xfkgtqxlj9

OpenStudy (anonymous):

that isnt working @jim_thompson5910

OpenStudy (anonymous):

i got B

OpenStudy (anonymous):

olve the first equation for y: y = (6 - 2x) substitute in equation 2... - 4x + 3(6 - 2x) = 8 - 4x + 18 - 6x = 8 - 10x = - 10 x = 1 1,4 correct @imqwerty ?

jimthompson5910 (jim_thompson5910):

do you see 2 lines when you click the link?

jimthompson5910 (jim_thompson5910):

you should see this